A 0.25kg ball is released from a height of 15m. The ball hits the floor and bounces in the opposite direction with 80% of its velocity. Calculate the impulse imparted by the floor on the ball.
1
Expert's answer
2020-07-06T15:40:49-0400

The velocity of the ball before collision with floor

"v_i=\sqrt{2gh}=\sqrt{2\times 9.8\times 15}=17.1\:\rm m/s"

The velocity of the ball after collision with floor

"v_f=0.8v_i=0.8\times 17.1=13.7\:\rm m/s"

Hence, the impulse imparted by the floor on the ball


"\Delta p=m|{\bf v}_f-{\bf v}_i|=m(v_f+v_i)\\
=0.25\times(13.7+17.1)=7.7\:\rm kg\cdot m/s"